OpenStudy (anonymous):

What is the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (anonymous):

Please help me!!!

OpenStudy (anonymous):

@FateRocks8D ?

OpenStudy (anonymous):

I'm like so lost

OpenStudy (anonymous):

xD me too!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

who is that?

jimthompson5910 (jim_thompson5910):

The slope of this parallel line is -3 since parallel lines have equal slopes. m = -3 (x,y) = (-2,4) y = mx+b y = -3x + b 4 = -3(-2) + b solve for b